OpenStudy (anonymous):

Michael drew a triangle with the sides measuring 12.5 centimeters, 30 centimeters, and 32.5 centimeters. Using the Pythagorean theorem, determine if Michael’s triangle is a right triangle. On the lines below, explain how you determined your answer.

OpenStudy (amistre64):

whats the pythag thrm state?

OpenStudy (anonymous):

okay, there is only so many ways to make a triangle using the formula: a^2+b^2=c^2. so now try all the diff ways and if one way works, then the answer is yes, it is possible.

OpenStudy (amistre64):

just try one way, since the longest side has to be the hypot

OpenStudy (anonymous):

the answer is yes:) 12.5^2+30^2=32.5^2

OpenStudy (anonymous):

The proportions are more obvious if you multiply by two. 25/60/65 is a 5/12/13 multiplied by five. (5/12/13 is an often-seen pythagorean triple used in textbook problems)